OpenStudy (anonymous):

I need a little bit of algebraic help, please :) I have already come to the conclusion and I'm sure it is correct but my answer doesn't match A, B, C, or D.

OpenStudy (azureilai):

whats the question?

OpenStudy (anonymous):

what are the answers

OpenStudy (anonymous):

Solve for Y in the first equation. x+3y=6 2x-y=10

OpenStudy (azureilai):

y= -1/3x +2 y=2x-10

OpenStudy (anonymous):

Can you add a Printscreen photo of the question & answers?

OpenStudy (anonymous):

Thanks for the help. My answer y=6... I'm not sure how I came to that but I thought it was right :P I should have read the question for carefully